# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ 6 Conclusion
In this work we have used the Worthey (1994) stellar population models to investigate stellar population age and abundance parameters and their radial gradients for 2417 SDSS MaNGA ETGs. The models are composite in metallicity with an ADF that resembles nearby galaxies. They also include an improvement, magnesium amplification, that allows them to approach true [X/H] and [X/R] values for the first time. Thus, these results supersede previous results by models without abundance-sensitive isochrone shifts, which, as of this writing, is all past work. That said, the improved results are not dramatically different, albeit with interesting new behavior in N and Na. The models are inverted to infer mean age, ADF peak [Fe/H], and mean elemental abundances for C, N, Na, and Mg for each elliptical annulus within each of the 2417 galaxies. We investigate primarily trends with $\sigma$ and [Fe/H] in this paper.
@@
In terms of galaxy formation and evolution, our findings are more consistent with formation of ETGs via merging (citations in §$\S$1) rather than strong monolithic collapse (citations in §$\S$1). Davies et al. (1993) proposes a scenario where smaller galactic subunits are formed by dissipational collapse which underwent dissipationless major mergers in subsequent times. This theory is supported in other works on stellar population parameter gradients (Kobayashi, 2004; Goddard et al., 2017). In our work, the strong dependence of [Fe/H] radial gradient values on $\sigma$ poses a problem to such theory. A merging dominant galactic evolution pathway is more apt in explaining our results. The radial gradients in C, N, and Mg also correspond well with the negative radial gradient in [Fe/H] pointing to the fact that overall element enrichment ($Z$) varies with galactocentric radius, with metal-rich centers and relatively metal-poor outer parts (Mehlert et al., 2003). While some cosmological simulations of galaxy formation support a two-phase formation model (Oser et al., 2010), our findings suggest that merger events were the primary drivers of galaxy evolution, with a minor contribution from dissipational collapse.
@@
The internal scatter ($\delta_{i}$) calculated in this work is an inference only rarely exploited in the literature. We found a decreasing trend for $\delta_{i}$ in both the age and [Fe/H] with the $\sigma$. The $\delta_{i}$ in radial gradients in those two quantities are sufficient to explain the scatter in the radial gradients calculated for them. The $\delta_{i}$ in elemental abundances and their gradients show in general lower $\delta_{i}$ values compared to the age and [Fe/H]. The only exception being Na. It is worth investing in the future the stand alone behaviour of Na compared to other elements considered in this work. As a note for future work, we also intend to study the effect of environment on the intrinsic scatter of astrophysical parameters.
Aside from substantial zero point and scale shifts in the abundances, we found a remarkable similarity between our work and the IllustrisTNG simulation in the abundance-$\sigma$ slope inflection that occurs at log $\sigma\approx 2$. The IllustrisTNG simulation result of [Fe/H] decreasing (for log($\sigma$)$>$2.0) with $\sigma$ matches very well with our work. Also, the light elements (C, N, and Mg) show a prominent increased abundance for log($\sigma$)$>$2.0 both in IllustrisTNG and our work. This potentially points to better and more efficient chemical enrichment in light elements in the subunits that eventually merge to form the largest ETGs.
@@ Acknowledgements
Appendix A Plots showing comparison with Illustris
This section of the appendix contains figures 28, 29, 30, 31, 32, and 33 elucidating comparison between our results and IllustrisTNG results for age, [Fe/H], [C/R], [N/R], [Mg/R], and ADF respectively.
@@ Appendix A Plots showing comparison with Illustris
Figure 28: The left panel shows how the mean age for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the Illustris and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the age of each individual Illustris galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 29: The left panel shows how the mean [Fe/H] for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the Illustris and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [Fe/H] of each individual Illustris galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 30: The left panel shows how the mean carbon abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the Illustris and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [C/R] of each individual Illustris galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 31: The left panel shows how the mean nitrogen abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the Illustris and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [N/R] of each individual Illustris galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 32: The left panel shows how the mean magnesium abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the Illustris and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [Mg/R] of each individual Illustris galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
Figure 33: The Abundance Distribution Function (ADF) for individual Illustris galaxies (in colored solid lines) plotted alongwith ADF used in this work (black lines with different styles). In this work, we have specifically used the ‘Normal’ ADF which resembles well with the Illustris ADF.
@@ Appendix B Spectral Feature Table
This section of the appendix contains a part of the table showing the details of the spectral features used in conjunction with the Worthey (1994) model to calculate age, [Fe/H], and elemental abundances. The full table is available online at [link to be inserted].
@@ Appendix C Monte Carlo (MC) Results
This section of the appendix shows a part of the table with the results from our MC simulation. The full table is available at [link to be inserted].
Table 7: Table showing the values of our derived astrophysical parameters, their radial gradients, and corresponding errors. This is a part of the table and the full table is available online at [link to be inserted].

### v2

@@ 6 Conclusion
We use updated Worthey (1994) stellar population models to investigate stellar population age and abundance parameters and their radial gradients for 2968 SDSS MaNGA ETGs. The models are composite in metallicity with an ADF that resembles nearby galaxies. They also include an improvement, magnesium amplification, that allows them to approach true [X/H] and [X/R] values for the first time. Thus, these results supersede previous results by models without abundance-sensitive isochrone shifts, which, as of this writing, is all past literature. That said, the improved results are not dramatically different, albeit with interesting new behavior in N and Na. The models are inverted to infer light-weighted mean age, ADF peak [Fe/H], and mean elemental abundances for C, N, Na, and Mg for each elliptical annulus within each of the 2968 galaxies. We investigate primarily trends with $\sigma$ and [Fe/H] in this paper.
@@
In terms of galaxy formation and evolution, our findings are more consistent with formation of ETGs via merging (citations in $\S$1) rather than strong monolithic collapse (citations in $\S$1). Davies et al. (1993) proposes a scenario where smaller galactic subunits are formed by dissipational collapse which underwent dissipationless major mergers in subsequent times. This theory is supported in other works on stellar population parameter gradients (Kobayashi, 2004; Goddard et al., 2017). In our work, the strong dependence of [Fe/H] radial gradient values on $\sigma$ poses a problem to such theory. A merging dominant galactic evolution pathway is more apt in explaining our results. The radial gradients in C, N, and Mg also correspond well with the negative radial gradient in [Fe/H] pointing to the fact that overall element enrichment ($Z$) varies with galactocentric radius, with metal-rich centers and relatively metal-poor outer parts (Mehlert et al., 2003). While some cosmological simulations of galaxy formation support a two-phase formation model (Oser et al., 2010), our findings suggest that merger events were the primary drivers of galaxy evolution, with a minor contribution from dissipational collapse.
@@
The astrophysical scatter ($\delta_{i}$) calculated in this work is an inference only rarely exploited in the literature. We found a decreasing trend for $\delta_{i}$ with $\sigma$ in both age and [Fe/H]. The $\delta_{i}$ in those two quantities are sufficient to explain the scatter in the radial gradients calculated for them. The $\delta_{i}$ in elemental abundances and their gradients show in general lower $\delta_{i}$ values compared to the age and [Fe/H]. The only exception to this rule is Na. For the future, it is worth investigating the standalone behavior of Na. Also for future work, we intend to study the effect of environment on the intrinsic scatter of astrophysical parameters.
Comparing with TNG simulations, we find substantial zero point and scale shifts in the ages, though we note that our population models do not yet include blue straggler stars. The TNG simulation result of [Fe/H] decreasing with $\sigma$ (for log($\sigma$)>>2.0) matches our results. Also, the light elements (C, N, and Mg) show an abundance slope inflection for log($\sigma$)>>2.0 both in TNG and our work. This potentially points to better and more efficient chemical enrichment in light elements in the subunits that eventually merge to form the largest ETGs.
@@ Acknowledgements
Appendix A Plots that compare with TNG
This section of the appendix contains figures 30, 31, 32, 33, 34, and 35 elucidating comparison between our results and TNG results for age, [Fe/H], [C/R], [N/R], [Mg/R], and ADF respectively.
@@ Appendix A Plots that compare with TNG
Figure 30: The left panel shows how the mean age for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the TNG and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the age of each individual TNG galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 31: The left panel shows how the mean [Fe/H] for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the TNG and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [Fe/H] of each individual TNG galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 32: The left panel shows how the mean carbon abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the TNG and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [C/R] of each individual TNG galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 33: The left panel shows how the mean nitrogen abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the TNG and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [N/R] of each individual TNG galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
@@
Figure 34: The left panel shows how the mean magnesium abundance for different $\sigma$ bins are changing with distance from the center of the galaxy. The solid lines show results from the TNG and the stars with errorbars show results from our work. Both the lines and symbols use same color codes for different $\sigma$ bins. The right panel shows how the [Mg/R] of each individual TNG galaxy (in red plus) and age from 3rd annulus of each MaNGA galaxy (in blue star) are changing with $\sigma$. MaNGA galaxies have errors associated with each data point.
Figure 35: The Abundance Distribution Function (ADF) for individual TNG galaxies (in colored solid lines) plotted along with three ADFs from an analytic function that is used in the Worthey CSPs. In this work, the ‘Normal’ ADF (dotted black) is used. compfit centers the ADF peak anew for each galaxy annulus, but for illustration purposes the model ADF peaks are drawn centered on zero.
@@ Appendix B Spectral Feature Table
This section of the appendix contains a part of the table showing the details of the spectral features used in conjunction with the Worthey (1994) model to calculate age, [Fe/H], and elemental abundances. The full table is available upon request to the authors.
@@ Appendix C Monte Carlo (MC) Results
This section of the appendix shows a part of the table with the results from our MC simulation. The full table is available upon request to the authors.
Table 7: Table showing the values of our derived astrophysical parameters, their radial gradients, and corresponding errors. This is a part of the table and the full table is available upon request to the authors.